OpenStudy (anonymous):

a=2t+Qa/2t+q solve for Q I keep getting stuck. if anyone can help thank you!

OpenStudy (anonymous):

is it \[a=\frac{2t+Qa}{2t+q}\]

OpenStudy (anonymous):

yes,just like that

OpenStudy (anonymous):

and it is the capital \(Q\) you are looking for right?

OpenStudy (anonymous):

yes, the Capital Q.

OpenStudy (anonymous):

start with \[a(2t+q)=2t+Qa\] then subtract \(2t\) and get \[a(2t+q)-2t=Qa\] and finally divide by \(a\)

OpenStudy (anonymous):

there are lots of ways to write the answer but be careful

OpenStudy (anonymous):

you could write \[\frac{a(2t+q)-2t}{a}=Q\]

OpenStudy (ajspeller):